What is the x-coordinate of the ordered pair solution of the system below?
x-3y=9
5x+2y=11
A) -3
B) -2
C) 2
D) 3

Answer by macston(5194)

x-3y=9
x=3y+9 Use to substitute for x.
5x+2y=11 Substitute for x from above.
5(3y+9)+2y=11
15y+45+2y=11
17y=-34
y=-2
x-3y=9
x-3(-2)=9
x+6=9
x=3
ANSWER: x=3 Choice (D.)